the electric field 0.4 m from a very long unifoem line of charge is 840 N/C.How much charge is contained in a 2.00 cm section of the line ?

Grade:12

2 Answers

vikas askiitian expert
13 years ago

let the charge density of line of charge is d..

now we have

                        electric field by line of charge = 2kd/r      ( k is electrostatic constant =9.109 )

at  0.4m electric field is 840

 

so,     2kd/0.4=840

           d=168/k=(1.86).(10)-8 columb/m

charge density is d so the charge contained by any small portion is d(x)

charge contained in 2cm=(1.86)(2)(10)-10

                                  =3.73(10)-10 columb

 

Raghunath Yadav Karike
13 years ago

Dear Raju,

In this case we have the formula i.e, E=(lambda)/(2pi(Eo)r) ------------------> 1

Determine the value of lambda that is linear charge density from above and apply it for the second case.

By 1 we get the value of lambda.

then to calculate the required charge use **** lambda=charge/length

Then u can easily get the answer